OpenStudy (anonymous):

Please help me with this problem. Find the volume lying the sphere x^2+y^2+z^2=2z and inside the paraboloid z=x^2+y^2

OpenStudy (experimentx):

x^2+y^2+z^2=2z << you sure??

OpenStudy (anonymous):

Yes please

OpenStudy (experimentx):

the sphere is degenerate ... has radius zero.

OpenStudy (anonymous):

ok

OpenStudy (tkhunny):

@Brotherman x^2 + y^2 + z^2 - 2z + 1 = 1 x^2+y^2+(z-1)^2 = 1 The sphere is not degenerate.

OpenStudy (amistre64):

i was thinking radius of 1 as well ...

OpenStudy (experimentx):

woops!! sorry my bad :((

OpenStudy (experimentx):

solving those two equations we get x^2+y^2+(z-1)^2 = 1 z=x^2+y^2 ---------------------- z = 0 to z = x^2+y^2 = 1